You are studying a population of beetles whose population size was initially 2000. After one month of observations, you record 300 births and 125 deaths. Calculate b, d and r for this population. Based on your calculations, how many beetles will there be in another month?
Answer #1

SOLUTION:

We know that there are 300 births in the population over the month, in our population of 2,000 individuals; we can express this as a rate by doing the following: (rate if the answer is helpful)

Birth rate (b) = 300/2000 = 0.15 births/(indiv. x month)

Using the same logic…

Death rate (d) = 125/2000 = 0.0625 deaths/(indiv. x month)

r = birth rate - death rate = 0.15-0.0625=0.0875

Nt= Noert

We know that t = 1 month (given in the question)

Therefore, Nt= 2000e(0.0875)(1)

Nt= 2000e0.0875

Taking ln on both sides (ln Nt=ln No+r)

ln Nt=ln2000+0.0875

ln Nt= 7.6+0.0875

ln Nt= 7.6875

Now (300-125=175)

175*7.6875=1345.3125

By adding this with initial population we will get

2000+1345.3125=3345.3125

=~3300 beetles or (3000- 3300)